OpenStudy (anonymous):

Suppose the derivative of f exists, and assume that f(1)=3, and f′(1)=5. Let g(x)=x^2 f(x), and h(x)=[f(x)]/[x−0]. a) g′(1)=.......................... Find the equation of the tangent line to g(x) at x=1. y=.......................... b) h′(1)=.......................... Find the equation of the tangent line to h(x) at x=1. y=........................... Hint: Find g′(x) and h′(x) in terms of f(x) and f′(x). You just need to use the basic product and quotient rules. The equation of a tangent line to a function k(x) at x=x0 is: y=k′(x0)(x−x0)+k(x0).

OpenStudy (campbell_st):

for g'(x) you'll need to product rule... do you know it..?

OpenStudy (anonymous):

product rule of x^2 = 2x?

OpenStudy (campbell_st):

so do you know the derivative of g'(x) ... and thats only part of the product rule.

OpenStudy (anonymous):

can you tell me the equation i am supposed to find derivative of?

OpenStudy (anonymous):

but, use the f(x) as your (x,y) (1,3) Hope this helps.

OpenStudy (campbell_st):

ok... for g(x) let u = x^2 du/dx = 2x v = f(x) du/dx = f'(x) so using the product rule \[g'(x) = u \times \frac{dv}{dx} + v \times \frac{du}{dx}\] so you will have \[g'(x) = x^2 \times f'(x) + 2x \times f(x)\] you need to find \[g'(1) = (1)^2 \times f'(1) + 2(1) \times f(1)\] f(1) and f'(1) are given above... substitute and evaluate which will give the gradient at x = 1 to find your point evaluate g(1) = (1)^2 x f(1) so you point will be (1 , g(1)) then find the equation.

OpenStudy (anonymous):

so g'(1) = 1 x 5 + 2 x 3=11

OpenStudy (campbell_st):

that correct... so thats the slope at x = 1 you now need to find g(1) ... g(1) = (1)^2 * f(1)

OpenStudy (anonymous):

g(1) = 3

OpenStudy (campbell_st):

ok... so you need to find the equation of a straight line( the tangent) with m = 11 and the point (1, 3) can you do that..?

OpenStudy (anonymous):

y-3=11(x-1)?

OpenStudy (anonymous):

33x-33

OpenStudy (campbell_st):

isn't it y - 3 = 11(x - 3) so y - 3 = 11x - 33 or y = 11x - 30

OpenStudy (campbell_st):

ready for h(x)?

OpenStudy (anonymous):

yup. all processed

OpenStudy (anonymous):

so should it be: f'(x)(x-0)-f(x)(x-0)/(x-0)^2?

OpenStudy (campbell_st):

ok... so for the quotient rule ok... almost correct... you just need the derivative of x - 0

OpenStudy (anonymous):

yes yes of course

OpenStudy (campbell_st):

so you will have \[h'(1) = \frac{ f'(1)\times (1 - 0) - f(1) \times 1}{(1 - 0)^2}\]

OpenStudy (anonymous):

so.....-10/1?

OpenStudy (anonymous):

=-10

OpenStudy (campbell_st):

ok... so thats the gradient... you need to find h(1) as the point for the line is (1, h(1))

OpenStudy (anonymous):

just a wild guess: will it be 3/(1-0)?

OpenStudy (anonymous):

=3?

OpenStudy (campbell_st):

ok.. so find the equation of the tangent where m = -10 and the point is (1, 3) can you do that..?

OpenStudy (anonymous):

y-3= -10(x-3)?

OpenStudy (anonymous):

=-30x=90?

OpenStudy (campbell_st):

ok... thats great

OpenStudy (campbell_st):

so y - 3 = -10(x - 1) y - 3 = -10x + 10

OpenStudy (campbell_st):

then y = -10x + 13 should be the equation of the tangent

OpenStudy (anonymous):

the computer shows that the only right answer i got out of the 4 is the first one.g'(1)=11

OpenStudy (anonymous):

y=11x-30 (incorrect) h'(1)= -10 (incorrect) y=-10x+13(incorrect)

OpenStudy (campbell_st):

ok... the tangent for the 1st line is y - 3 = 11(x - 1) so y -3= 11x - 11 y = 11x - 8 try than for the tangent to g(x)

OpenStudy (anonymous):

you got it right

OpenStudy (campbell_st):

and for h(x) you should have \[h'(1) = \frac{( 1- 0) time 5 - (3 \times 1)}{1} = 2\] so the value of h'{1} = 2 and y - 3 = 2( x - 1) y - 3 = 2x - 2 y = 2x + 1 is the tangent

OpenStudy (anonymous):

YES! YES! YES!

OpenStudy (anonymous):

thank you very much campbell

OpenStudy (campbell_st):

ok... just arithmetic errors... hope it all helped.

OpenStudy (anonymous):

yes, very much

OpenStudy (anonymous):

you want to work on more questions maybe?

OpenStudy (campbell_st):

nope... I think you need to go and practice the product and quotient rules...

OpenStudy (anonymous):

ok great. thanks again and see you around